OpenStudy (anonymous):

Write and Solve inequality On A trip from Virginia to Florida, the Sampson family wants to travel at least 420 miles in 8 hours of driving. What must be their average rate of speed? explain how ya got ur answer Write and Solve inequality On A trip from Virginia to Florida, the Sampson family wants to travel at least 420 miles in 8 hours of driving. What must be their average rate of speed? explain how ya got ur answer @Mathematics

OpenStudy (anonymous):

since distance equals rate times time, and since you know both the time (8 hours) and the distance (420 miles) we can call the rate say \(r\) and write \[8r\leq 420\]

OpenStudy (anonymous):

whoa is that wrong! "at least" try \[8r\geq 420\]

OpenStudy (anonymous):

so their average rate of speed must be at least \[r\geq 420\div 8=52.5\] in other words at least 52.5 mph